OpenStudy (anonymous):

HELP? Which of the following is the simplified form of x plus 5 over x squared minus 25 times the fraction x minus 5 over x squared minus 10 x plus 25 ? 1 over the quantity x minus 5 end quantity squared 1 over the quantity x plus 5 end quantity squared 1 over x squared minus 25 1 over x squared plus 25

OpenStudy (rmrjr22):

x plus 5 = (x+5) over = / x squared minus 25 times the fraction x minus 5 over x squared minus 10 x plus 25 x^2 - 25((x-5)/(x^2 - 10x+25)

OpenStudy (anonymous):

huh?

OpenStudy (rmrjr22):

\[(x+5)/x^2 - 25((x-5)/(x^2 - 10x+25))\]

OpenStudy (rmrjr22):

unfactor everything until you get what ever is on the top. In this case ...(x+5)

OpenStudy (rmrjr22):

x^2−25((x−5)= (x+5)(x-5)(x-5)

OpenStudy (anonymous):

ok...

OpenStudy (rmrjr22):

x^2−10x+25 = (x-5)(x-5)

OpenStudy (rmrjr22):

now just cancel out top from bottom. Any like terms like (x-5) or (x+5)

OpenStudy (anonymous):

OK?

OpenStudy (anonymous):

I got D.

OpenStudy (rmrjr22):

yes

OpenStudy (anonymous):

That's right?

OpenStudy (rmrjr22):

believe so

OpenStudy (anonymous):

Alright, thanks.

OpenStudy (rmrjr22):

I actually ended up with 1...

OpenStudy (luigi0210):

Hm.. Kind of hard to see without the equation editor You sure?

OpenStudy (anonymous):

>.<

OpenStudy (anonymous):

Oh, you might be right then...

OpenStudy (rmrjr22):

Did on pen and paper...

OpenStudy (luigi0210):

@rmrjr22 what was your final answer?

OpenStudy (rmrjr22):

1...

OpenStudy (anonymous):

Like number 1?

OpenStudy (rmrjr22):

yes... everything cancels out

OpenStudy (anonymous):

then you are probably right, so A it is.

OpenStudy (luigi0210):

Okay wait show me your equation

OpenStudy (rmrjr22):

mine?

OpenStudy (luigi0210):

Yes

OpenStudy (anonymous):

\[\frac{ x+5 }{ x^2-25 }\times \frac{ x-5 }{ x^2-10+25 }\]

OpenStudy (anonymous):

Oh.. dangit.

OpenStudy (rmrjr22):

\[(x+5)/x^2−25((x−5)/(x^2−10x+25)) = x+5/((x-5)(x-5))((x-5)/(x-5)(x-5))\]

OpenStudy (luigi0210):

@dreamer05 with that equation we end up with \[\frac{ 1 }{ (x-5)^2 }\]

OpenStudy (luigi0210):

So it was a

OpenStudy (anonymous):

You sure??

OpenStudy (rmrjr22):

yup

OpenStudy (luigi0210):

Haha, sorry about that

OpenStudy (luigi0210):

@rmrjr22 you should try using the fraction function.. it'll make things easier to read

OpenStudy (rmrjr22):

haha yeah noticed

OpenStudy (anonymous):

cus one was \[\frac{ 1 }{ x^2-10+25 }\]

OpenStudy (luigi0210):

Good work guys

OpenStudy (anonymous):

So it is A.? Cu that;s A answer.

OpenStudy (luigi0210):

Yup, it's A

OpenStudy (anonymous):

Okay fanks c: